OpenStudy (anonymous):

Find the perimeter and area of a rectangle with a length of 3.2 m and a width of 1.5 m. P = 4.7 m; A = 2.25 m2 P = 4.7 m; A = 4.8 m2 P = 9.4 m; A = 2.4 m2 P = 9.4 m; A = 4.8 m2

OpenStudy (anonymous):

@pooja195

Mehek (mehek14):

\(\tt{P=3.2*2+1.5*2\\A=3.2*1.5}\)

OpenStudy (anonymous):

4.8

Mehek (mehek14):

which one?

OpenStudy (anonymous):

d

Mehek (mehek14):

Yes that's correct :]

OpenStudy (anonymous):

thanks!!

Mehek (mehek14):

yw ^_^